# Topological advantage for adsorbate chemisorption on conjugated chains

### v2

@@ I Introduction
Topological insulators are electronic phases of matter with a bulk band gap and gapless surface states protected by global symmetries [hasan2010colloquium, moore2010birth, schnyder2008classification, asboth2016short, kane2005quantum, kane2005z, bernevig2006quantum]. These materials exhibit intriguing properties, including high surface carrier mobility [shekhar2012ultrahigh, shekhar2012electronic, analytis2010two, thio1998giant, butch2010strong, jia2011low, li2012electron] with low power dissipation [roth2009nonlocal, chang2015zero], and spin-polarized currents with unconventional textures [ando2014electrical, tian2015electrical, souma2011direct, pan2011electronic]. Potential applications include spintronics [hsieh2009observation, he2019topological, fan2016spintronics, he2022topological, yokoyama2014spintronics], quantum computing [he2019topological, paudel2013three, scappucci2021crystalline], and thermoelectrics [xu2017topological, ivanov2018thermoelectric, matsushita2017thermoelectric, muchler2013topological, xu2016thermoelectric]. Recently, there has also been interest in utilizing topological insulators and semimetals in electrochemistry and chemical catalysis [rajamathi2017weyl, schoop2018chemical, li2020heterogeneous, muechler2020topological, narang2021topology, zhang2023topological, weng2024understanding, li2023topological, li2024synergistic, lau2025facet, laderer2025topological]. Their highly localized conducting boundary modes may significantly influence surface reactions, and their properties are protected against local perturbations (e.g., lattice defects and impurities). For photocatalysis, high surface mobility is desirable because efficient charge separation and suppressed electron–hole recombination can enhance performance [zhang2016mnpse3, lv2017two, li2020heterogeneous, rouzhahong2020first]. Indeed, recent experimental [chen2013study, rajamathi2017weyl, li2019surface, he2019topological1, yang2019electrochemical] and computational [chen2011co, Li2017, weng2024understanding] studies provide evidence that, under certain conditions, topological matter exhibits enhanced reactivity mediated by topologically protected boundary electrons. Nevertheless, the extent to which the surfaces of topological materials can be systematically exploited for high-efficiency, selective synthesis remains an open question.
@@
In heterogeneous catalysis, adsorbed molecules undergo several processes on a solid surface [hoffmann1988chemical, ertl1990elementary, jiang2019dynamics, second2008handbook]. A central step is charge transfer between the extended material and the adsorbates [weisz1953effects, lewis2005chemical, wodtke2008energy]. Orbital hybridization and electron donation into a molecular system influence adsorption [sung1985carbon, hoffmann1988chemical, santen1991theoretical, jung2021understanding], surface diffusion [mckenna2010interplay, wahlstrom2004electron], and desorption, and can reduce energetic barriers for bond breaking or association [kasemo1996charge, dai1995laser]. Likewise, adsorbate–surface energy exchange via molecular vibrational relaxation and excitation of boundary electron–hole pairs (EHPs) [newns1969self, bohnen1975friction, persson1980vibrational, hellsing1984electronic, tully1990molecular, tully1995molecular, tully2000chemical, alec2004electronically, dou2020nonadiabatic, ghan2023interpreting] has been implicated in $\text{H}_{2}$ relaxation on metal surfaces [persson1982electronic, juaristi2008role], $\text{CO}_{2}$ reduction [nam2020molecular, schneider2012thermodynamics, diercks2018role, ge2018electron], and water oxidation [materna2017anchoring]. Here we focus on two mechanisms that are important to heterogeneous catalysis and can be directly analyzed in a minimal topological setting: (i) charge hybridization between molecular frontier orbitals and boundary electronic states, and (ii) nonadiabatic vibrational energy dissipation (electronic friction) mediated by EHPs. We ask how these processes are modified when the substrate hosts symmetry-protected boundary modes versus trivial or metallic electronic structures.
We employ the Su–Schrieffer–Heeger (SSH) model [su1979solitons, heeger1988solitons] as a minimal platform that supports topological, metallic, and trivial insulating regimes within a common Hamiltonian and enables controlled introduction of defects and detailed analysis in the thermodynamic limit. While the SSH chain is one-dimensional and spinless it captures the essential physics of symmetry-protected boundary modes that underlie many proposed catalytic advantages. Our goal is to isolate the role of topology from complexities of specific materials that may hinder a clear interpretation of the observed effects of topological materials on adsorbates. We study prototype molecules coupled to the chain and analyze the electron occupancy of the lowest unoccupied molecular orbital (LUMO) for single and multiple adsorbates, both with and without trivial and topological defects. We also report how crossing the topological phase transition manifests in the electronic friction experienced by an adsorbate, computed via a linear-response kernel that quantifies EHP-mediated vibrational damping at the molecule–substrate contact (see Sec. II). Within the parameter ranges explored, we find a consistent advantage of the topological phase over the metallic phase for charge hybridization, and a significant enhancement of electronic friction in the topological phase relative to the trivial insulator.
This article is organized as follows. In Sec. II, we define the effective Hamiltonian for the adsorbate–chain system and introduce the molecular observables and configurations investigated. In Sec. IIIA, we present observations of topologically enhanced chemisorption. In Sec. IIIB, we analyze the mean LUMO electron occupancy in the presence of variable concentrations of molecules and defects. Section IIIC discusses electronic friction across the topological, trivial, and metallic phases. We summarize the main results and outline future directions in Sec. IV. Appendix A details the numerical method used to compute electronic friction.
II Theory
II.1 Fano-Anderson SSH model
@@
The simplest scenario we considered consists of a molecule adsorbed at different regions of an open conjugated SSH chain (Fig. 1). The isolated molecule has a closed shell and a single low-lying empty electronic orbital, while the SSH chain models the extended material [su1979solitons, su1980soliton]. The molecule interacts with nearby lattice sites as in the Fano–Anderson model [fano1961effects, anderson1961localized]. To probe the roles of bulk versus boundary modes, we place the molecule either near the chain edge or at the center (Fig. 1). f The SSH model represents a polyacetylene chain by a 1D tight-binding lattice with two identical sites per unit cell, nearest-neighbor staggered hopping [su1979solitons], and one electron per site. For a spinless open system, the isolated chain Hamiltonian is
@@
$\quad+w\sum_{j=1}^{N-1}\!\left(\hat{d}_{j+1,A}^{\dagger}\hat{d}_{j,B}+\hat{d}_{j,B}^{\dagger}\hat{d}_{j+1,A}\right),$ (1)
where $v$ ($w$) is the intracell (intercell) hopping, $N$ is the number of unit cells, and $\hat{d}_{j,\alpha}^{\dagger}$ ($\hat{d}_{j,\alpha}$) creates (annihilates) an electron on sublattice $\alpha\in\{A,B\}$ in cell $j$. The gapped phases correspond to $v\neq w$: $v>w$ is a trivial insulator without boundary-localized modes, whereas $v<w$ yields a gapped system with two zero-energy boundary modes [su1979solitons, asboth2016short].
@@
$\quad+T_{N/2-1,B}\,\hat{d}_{N/2-1,B}^{\dagger}\hat{d}_{0}\;+\;\text{h.c.},$ (4)
@@
Figure 1: Schematic Fano–Anderson SSH setup. White and orange circles denote sublattices A and B, respectively, each with a single $p_{z}$ orbital. Intra- and intercell hoppings ($v$ and $w$) are tunable. A diatomic adsorbate couples either near an edge ($x_{M}=x_{E}$) or at the chain center ($x_{M}=x_{B}$). (a) Hybridization with edge/bulk modes in the topological phase ($v<w$). (b) Hybridization with a trivial insulator ($v>w$) or with the metallic limit ($v=w$), where states are delocalized.
Polyacetylene chains may host both trivial and topological lattice defects [su1979solitons, heeger1988solitons]. The latter are domain walls separating trivial ($v>w$) and topological ($v<w$) regions and support mid-gap localized modes. We therefore also consider a scenario where multiple molecules interacting with such defects. In the considered static limit (immobile defects), $N_{d}$ solitons centered at sites $m_{1},\dots,m_{N_{d}}$ are modeled by site-dependent nearest-neighbor hoppings
@@
with [takayama1980continuum]
@@
where $\xi$ is the soliton width. In disordered samples, each adsorbate is placed at a soliton center $m_{i}$ and couples to the nearest site with strength $T_{m_{i}}$ and to its two nearest neighbors $m_{i}\!\pm\!1$ with $T_{m_{i}}/3$, so each adsorbate interacts with three sites. The adsorbates are typically sufficiently far from each other that the adsorbate ensemble Hamiltonian is the the straightforward generalization of $\hat{H}_{0}$ to multiple noninteracting adsorbates.
II.2 Observables
To quantify charge transfer and hybridization for a single adsorbate, we compute the adsorbate population $n_{0}(R,x_{M})$ for a half-filled SSH chain at zero temperature with the Fermi energy set to $\mu=0$. Denoting by $\hat{P}_{\mathrm{occ}}(R,x_{M})=\Theta\!\big(\mu-\hat{H}(R,x_{M})\big)$ the projector onto the occupied single-particle subspace, where $\Theta$ is the Heaviside step function,
@@ II.2 Observables
To probe vibrational energy exchange with the substrate, we consider the adsorbate vibrational degree of freedom $R$ as a dynamical variable, modeled as a harmonic oscillator with mass $m$, frequency $\omega$, and displacement from equilibrium (at the electronic ground-state) $R$. We assume a linear vibronic coupling of the adsorbate level [persson1982electronic, dou2020nonadiabatic],
@@
$\varepsilon_{0}(R)=\varepsilon_{d}+\kappa R,\quad U_{0}(R)=\tfrac{1}{2}m\omega^{2}R^{2},$ (10)
where $\varepsilon_{d}$ is the LUMO energy at $R=0$, $\kappa=\partial\varepsilon_{0}/\partial R$ is the (Condon) linear vibronic coupling, and the dependence on $x_{M}$ is omitted for notational simplicity.
In contact with an extended electronic system, vibrational relaxation can proceed via nonadiabatic excitation of electron–hole pairs (EHPs) [persson1980vibrational, dou2020nonadiabatic]. Under the standard separation of timescales where the electronic subsystem equilibrates much faster than the nuclear motion [spohn1980kinetic, dann2018time], the vibrational energy loss rate is captured by an electronic friction kernel [bohnen1975friction, persson1980vibrational, tully1995molecular, huang2000vibrational, bunermann2015electron, ryabinkin2017mixed, dou2020nonadiabatic]. For a single coordinate $R$,
@@
where the trace is over the single-particle electronic space and $\hat{P}(\epsilon;R)=\sum_{\lambda}\delta\!\big(\epsilon-\epsilon_{\lambda}(R)\big)\,|\lambda(R)\rangle\langle\lambda(R)|$ is the spectral projector. For the Hamiltonian in (2) with (9), assuming (i) only the adsorbate level $\varepsilon_{0}(R)$ depends on $R$ (no $R$-dependence in the couplings $T_{s,\alpha}$: non-Condon effects neglected), and (ii) linear response about a fixed $R$, the friction reduces to
$\gamma(R)$ $=-\pi\hbar\,\kappa^{2}\!\int_{-\infty}^{\infty}\!\mathrm{d}\epsilon\sum_{\lambda,\lambda^{\prime}}\big|\langle\lambda(R)\,|\,\varepsilon_{0}\rangle\big|^{2}\delta\!\big[\epsilon-\epsilon_{\lambda}(R)\big]$
@@
The Dirac delta function is represented by a Gaussian function
@@
with a width $\sigma$ chosen on the order of half the HOMO–LUMO spacing of the finite open SSH chain near the metallic limit ($v\approx w$). This choice yields the smooth friction coefficients reported in Sec. III.3.
III Results and Discussion
III.1 Bulk and boundary chemisorption
In this section, we examine the chemisorption of a single adsorbate on the SSH chain under various conditions. Figures 2(a,b) report the LUMO occupation $n_{0}(R,x_{M})$ at 0 K and half filling ($\mu=0^{-}$) for adsorption near the edge ($x_{M}=x_{E}$) or at the bulk center ($x_{M}=x_{B}$) as a function of the dimerization ratio
@@ III.1 Bulk and boundary chemisorption
covering the trivial ($r<1$), metallic ($r\simeq 1$), and topological ($r>1$) regimes. Throughout this subsection the molecular geometry is fixed ($R=0$), so the adsorbate is characterized by its low-lying level energy $\varepsilon_{0}$ and placement $x_{M}$.
Figure 2(a) shows that the trivial insulator exhibits negligible electron sharing at the edge, whereas the topological phase displays strong hybridization with a sharp rise of $n_{0}(R,x_{E})$ upon entering $r>1$. The enhancement is most pronounced when $\varepsilon_{0}$ is nearly resonant with the boundary mode. Conversely, Fig. 2(b) shows that for adsorption in the bulk, significant $n_{0}(R,x_{B})$ occurs only in a narrow neighborhood of the critical point ($r\approx 1$), where the density of occupied states near $\varepsilon_{0}$ is largest. While the metallic phase maximizes bulk adsorption, boundary chemisorption in the topological phase remains the optimal overall scenario: $\max n_{0}(R,x_{B})<\max n_{0}(R,x_{E})=0.5$. Notably, appreciable adsorbate-edge hybridization persists even when $\varepsilon_{0}\gg 0$.
@@
The robustness of the topological enhancement is highlighted in Fig. 3, which directly compares edge adsorption in the topological phase to bulk adsorption in the metallic phase over varying $\varepsilon_{0}$ and couplings $T_{m,A}$. Across all cases examined, hybridization with metallic bulk states yields a lower $n_{0}$ than coupling to the topological edge mode, despite the larger metallic DOS near $\varepsilon_{0}$. Importantly, edge $n_{0}(R,x_{E})$ remains substantial even when $T_{1,A}\ll v,w$.
@@
Discussion. In the topological phase and near resonance ($\varepsilon_{0}\approx 0$), the adsorbate couples predominantly to the nearest edge mode. For $T_{1,A},T_{1,B}\ll v,w$ and $w/v-1\gg 0$, the hybridized edge–LUMO doublet is well isolated from bulk continua, yielding the symmetric occupancy $n_{0}(R,x_{E})\approx 0.5$. In the metallic regime ($r\simeq 1$), $\varepsilon_{0}\approx 0$ is near-degenerate with many occupied/unoccupied extended states and the resulting Fano line shape [fano1961effects] in the local molecular DOS produces a lower maximal $n_{0}\approx 0.4$. Off resonance, bulk adsorption admits the standard second-order estimate
$n_{0}(R,x_{B})\;\approx\;\sum_{E_{\kappa}^{(0)}<\mu}\frac{\bigl|\bm{T}_{c}\!\cdot\!\bm{u}_{\kappa,c}\bigr|^{2}}{\bigl(\varepsilon_{0}-E_{\kappa}^{(0)}\bigr)^{2}},$ (15)
where $E_{\kappa}^{(0)}$ denote bare SSH eigenenergies, $c=N/2$ and $\bm{T}_{c}$ and $\bm{u}_{\kappa,c}$ are
$\bm{T}_{c}\equiv\begin{pmatrix}T_{c,A}\\[2.0pt] T_{c,B}\\[2.0pt] T_{c-1,B}\end{pmatrix},\qquad\bm{u}_{\kappa,c}\equiv\begin{pmatrix}a_{c,\kappa}\\[2.0pt] b_{c,\kappa}\\[2.0pt] b_{c-1,\kappa}\end{pmatrix}$,
and $a_{j,\kappa},b_{j,\kappa}$ are the amplitudes of eigenstate $\kappa$ on sublattices $A/B$ at cell $j$. The sum runs over occupied states at T=0T\!=\!0 K. Away from criticality the occupied bulk states with non-negligible weight at the molecule are delocalized and detuned from $\varepsilon_{0}$, suppressing $n_{0}(R,x_{B})$. Conversely, when near $r\approx 1$ the increased density of states and small denominators in (15) enhance $n_{0}$.

### v3

# font=sf,footnotesize{tocentry}

@@ Abstract
Topological insulators are electronic phases of matter with a bulk band gap and gapless surface states protected by global symmetries hasan2010colloquium, moore2010birth, schnyder2008classification, asboth2016short, kane2005quantum, kane2005z, bernevig2006quantum. These materials exhibit intriguing properties, including high surface carrier mobility shekhar2012ultrahigh, shekhar2012electronic, analytis2010two, thio1998giant, butch2010strong, jia2011low, li2012electron with low power dissipation roth2009nonlocal, chang2015zero, and spin-polarized currents with unconventional textures ando2014electrical, tian2015electrical, souma2011direct, pan2011electronic. Potential applications include spintronics hsieh2009observation, he2019topological, fan2016spintronics, he2022topological, yokoyama2014spintronics, quantum computing he2019topological, paudel2013three, scappucci2021crystalline, and thermoelectrics xu2017topological, ivanov2018thermoelectric, matsushita2017thermoelectric, muchler2013topological, xu2016thermoelectric. Recently, there has also been interest in utilizing topological insulators and semimetals in electrochemistry and chemical catalysis rajamathi2017weyl, schoop2018chemical, li2020heterogeneous, muechler2020topological, narang2021topology, zhang2023topological, weng2024understanding, li2023topological, li2024synergistic, lau2025facet, laderer2025topological. Their highly localized conducting boundary modes may significantly influence surface reactions, and their properties are protected against local perturbations (e.g., lattice defects and impurities). For photocatalysis, high surface mobility is desirable because efficient charge separation and suppressed electron–hole recombination can enhance performance zhang2016mnpse3, lv2017two, li2020heterogeneous, rouzhahong2020first. Indeed, experimental chen2013study, rajamathi2017weyl, li2019surface, he2019topological1, yang2019electrochemical and computational chen2011co, Li2017, weng2024understanding studies provide evidence that, under certain conditions, topological matter exhibits enhanced reactivity mediated by topologically protected boundary electrons. Nevertheless, the extent to which the surfaces of topological materials can be systematically exploited for high-efficiency, selective synthesis remains an open question.
@@
In heterogeneous catalysis, adsorbed molecules undergo several processes on a solid surface hoffmann1988chemical, ertl1990elementary, jiang2019dynamics, ertl1997handbook. A central step is charge transfer between the extended material and the adsorbates weisz1953effects, lewis2005chemical, wodtke2008energy. Orbital hybridization and electron donation into a molecular system influence adsorption sung1985carbon, hoffmann1988chemical, santen1991theoretical, jung2021understanding, surface diffusion mckenna2010interplay, wahlstrom2004electron, and desorption, and can reduce energetic barriers for bond breaking or association kasemo1996charge, dai1995laser. Likewise, adsorbate–surface energy exchange via vibrational relaxation and excitation of boundary electron–hole pairs (EHPs) newns1969self, bohnen1975friction, persson1980vibrational, hellsing1984electronic, tully1990molecular, tully1995molecular, tully2000chemical, alec2004electronically, dou2020nonadiabatic, ghan2023interpreting, meng2025dynamics, preston2025nonadiabatic has been implicated in $\text{H}_{2}$ relaxation on metal surfaces persson1982electronic, juaristi2008role, $\text{CO}_{2}$ reduction nam2020molecular, schneider2012thermodynamics, diercks2018role, ge2018electron, and water oxidation materna2017anchoring. Here we focus on two mechanisms that are important to heterogeneous catalysis and can be directly analyzed in a minimal topological setting: (i) charge hybridization between molecular frontier orbitals and boundary electronic states, and (ii) nonadiabatic vibrational energy dissipation (electronic friction) mediated by EHPs. We ask how these processes are modified when the substrate hosts symmetry-protected boundary modes versus trivial or metallic electronic structures.
We employ the Su–Schrieffer–Heeger (SSH) model su1979solitons, heeger1988solitons as a minimal platform in which topological, metallic, and trivial insulating regimes arise within a single Hamiltonian, while still allowing controlled introduction of defects and systematic thermodynamic-limit analysis. Although the SSH chain is one-dimensional and we use a spinless, effective single-particle form, it faithfully captures the chiral (sublattice) symmetry and the associated boundary- and domain-wall-localized midgap orbitals that constitute the central ingredient of this work and that, as we show below, strongly modulate molecule–substrate interfacial properties.
This effective description is particularly well motivated for light-element $\pi$-conjugated polymers such as trans-polyacetylene: intrinsic spin–orbit coupling in the carbon $p_{z}$ manifold is weak compared with eV-scale hoppings boehme2013challenges, and electron–electron interactions bendazzoli1999density can be treated at an effective level through renormalized hopping amplitudes and onsite energies within a tight-binding picture. We briefly comment at the end on extensions to spinful models with strong spin–orbit coupling and/or strong correlations. The resulting minimal-model approach is therefore designed to isolate robust, design-relevant trends, and in particular to clarify when and where enhanced hybridization and nonadiabatic dissipation should be expected, in a way that enables a direct interpretation of the effects reported here.
Within this controlled setting, we couple molecular adsorbates to SSH-class substrates to isolate and quantify how symmetry-protected midgap states (at finite edges and at topological domain walls) govern molecule-substrate interfacial observables. Across metallic, trivial insulating, and topological regimes, and in the presence of both trivial and topological defects, we compute adsorbate level broadening (hybridization), LUMO occupancy, and charge transfer for single and multiple adsorbates, thereby disentangling the roles of the density of states and localized midgap modes. We further evaluate the electronic friction acting on an adsorbate within linear response, via a kernel that captures electron–hole-pair–mediated vibrational damping at the contact. We show that localized midgap modes can (i) strongly enhance adsorbate hybridization and charge donation relative to both trivial insulating and metallic substrates, and (ii) sharply increase vibrational energy dissipation through enhanced electronic friction upon crossing the topological transition; moreover, topological domain walls mitigate the potential measure-zero limitation of edge modes by providing a potentially extensive, chemically accessible set of such midgap states.
@@
Fano-Anderson SSH model. The simplest scenario we considered consists of a molecule adsorbed at different regions of an open conjugated SSH chain (Fig. 1). The isolated molecule has a closed shell and a single low-lying empty electronic orbital, while the SSH chain models the extended material su1979solitons, su1980soliton. The molecule interacts with nearby lattice sites as in the Fano–Anderson model fano1961effects, anderson1961localized. To probe the roles of bulk versus boundary modes, we place the molecule either near the chain edge or at the center (Fig. 1).
The SSH model represents a polyacetylene chain by a 1D tight-binding lattice with two identical sites per unit cell, nearest-neighbor staggered hopping su1979solitons, and one electron per site. For a spinless open system, the isolated chain Hamiltonian is
@@
$\qquad+w\sum_{j=1}^{N-1}\!\left(\hat{d}_{j+1,A}^{\dagger}\hat{d}_{j,B}+\hat{d}_{j,B}^{\dagger}\hat{d}_{j+1,A}\right),$ (1)
where $v$ ($w$) is the intracell (intercell) hopping, $N$ is the number of unit cells, and $\hat{d}_{j,\alpha}^{\dagger}$ ($\hat{d}_{j,\alpha}$) creates (annihilates) an electron on sublattice $\alpha\in\{A,B\}$ in cell $j$. The gapped phases correspond to $v\neq w$: $v>w$ is a trivial insulator without boundary-localized modes, whereas $v<w$ yields a gapped system with two zero-energy boundary modes su1979solitons, asboth2016short.
In a fully spinful description of a dimerized $\pi$-conjugated chain (e.g., polyacetylene), each spatially localized midgap bound state corresponds to a localized $\pi$ frontier orbital. Depending on filling and interactions, this orbital can be singly occupied and carry an unpaired electron, i.e., it acts as a radical center in the language of conventional organic chemistry bredas1984role. In the present work, we use a spinless SSH Hamiltonian, which can be viewed as isolating a single spin channel. In a more realistic spinful model the same spatial orbital would be twofold spin-degenerate. Hence, the topological midgap states of the SSH chain are tight-binding analogues of chemically familiar radical-like frontier orbitals, with the dimerization ratio ($w>v$ vs. $w<v$) controlling whether such orbitals are present or absent. In what follows, we focus on how the existence, localization, and filling of these midgap frontier orbitals modulate charge transfer and electronic friction between the conjugated backbone and the adsorbate.
@@
$\qquad+T_{N/2-1,B}\,\hat{d}_{N/2-1,B}^{\dagger}\hat{d}_{0}\;+\;\text{h.c.},$ (4)
@@
Figure 1: Schematic Fano–Anderson SSH setup. White and orange circles denote sublattices A and B, respectively, each with a single $p_{z}$ orbital. Intra- and intercell hoppings ($v$ and $w$) are tunable. A representative diatomic adsorbate couples either near an edge ($x_{M}=x_{E}$) or at the chain center ($x_{M}=x_{B}$). (a) Hybridization with edge/bulk modes in the topological phase ($v<w$). (b) Hybridization with a trivial insulator ($v>w$) or with the metallic limit ($v=w$), where states are delocalized.
Polyacetylene chains may host both trivial and topological lattice defects su1979solitons, heeger1988solitons. The latter are domain walls separating trivial ($v>w$) and topological ($v<w$) regions and support mid-gap localized modes. We therefore also consider a scenario where multiple molecules interacting with such defects. In the considered static limit (immobile defects), $N_{d}$ solitons centered at sites $m_{1},\dots,m_{N_{d}}$ are modeled by site-dependent nearest-neighbor hoppings
@@
with takayama1980continuum
@@
where $\xi$ is the soliton width. Although solitons in real polyacetylene generally arise as thermally activated defects, here we treat them as fixed static inhomogeneities and work at half filling with a fixed electron number. This approximation allows us to examine the fundamental role of topological defects without introducing additional assumptions regarding thermally driven soliton formation and dynamics.
In disordered samples, each adsorbate is placed at a soliton center $m_{i}$ and couples to the nearest site with strength $T_{m_{i}}$ and to its two nearest neighbors $m_{i}\!\pm\!1$ with $T_{m_{i}}/3$, so each adsorbate interacts with three sites. The adsorbates are typically sufficiently far from each other that the adsorbate ensemble Hamiltonian is the the straightforward generalization of $\hat{H}_{0}$ to multiple noninteracting adsorbates.
Observables. To quantify charge transfer and hybridization for a single adsorbate, we compute the adsorbate population $n_{0}(R,x_{M})$ for a half-filled SSH chain at zero temperature with the Fermi energy set to $\mu=0$. Defining the occupied single-particle projector as $\hat{P}_{\mathrm{occ}}(R,x_{M})=\Theta\!\big(\mu-\hat{H}(R,x_{M})\big)$, where $\Theta$ is the Heaviside step function, we write
@@
To probe vibrational energy exchange with the substrate, we consider the adsorbate vibrational degree of freedom $R$ as a dynamical variable, modeled as a harmonic oscillator with mass $m$, frequency $\omega$, and displacement from equilibrium (at the electronic ground-state) $R$. We assume a linear vibronic coupling of the adsorbate level persson1982electronic, dou2020nonadiabatic,
@@
$\varepsilon_{0}(R)=\varepsilon_{d}+gR,\qquad U_{0}(R)=\tfrac{1}{2}m\omega^{2}R^{2},$ (10)
where $\varepsilon_{d}$ is the LUMO energy at $R=0$, $g=\partial\varepsilon_{0}/\partial R$ is the (Condon) linear vibronic coupling, and the dependence on $x_{M}$ is omitted for notational simplicity.
In contact with an extended electronic system, vibrational relaxation can proceed via nonadiabatic excitation of electron–hole pairs (EHPs) persson1980vibrational, dou2020nonadiabatic. Under the standard separation of timescales where the electronic subsystem equilibrates much faster than the nuclear motion spohn1980kinetic, dann2018time, the vibrational energy loss rate is captured by an electronic friction kernel bohnen1975friction, persson1980vibrational, tully1995molecular, huang2000vibrational, bunermann2015electron, ryabinkin2017mixed, dou2020nonadiabatic. For a single coordinate $R$,
@@
where the trace is over the single-particle electronic space and $\hat{P}(\epsilon;R)=\sum_{\lambda}\delta\!\big(\epsilon-\epsilon_{\lambda}(R)\big)\,|\lambda(R)\rangle\langle\lambda(R)|$ is the spectral projector. For the Hamiltonian in (2) with (9), assuming (i) only the adsorbate level $\varepsilon_{0}(R)$ depends on $R$ (no $R$-dependence in the couplings $T_{s,\alpha}$, i.e., with non-Condon effects neglected), and (ii) linear response about a fixed $R$, the friction reduces to
$\gamma(R)$ $=-\pi\hbar\,g^{2}\!\int_{-\infty}^{\infty}\!\mathrm{d}\epsilon\sum_{\lambda,\lambda^{\prime}}\big|\langle\lambda(R)\,|\,\varepsilon_{0}\rangle\big|^{2}\delta\!\big[\epsilon-\epsilon_{\lambda}(R)\big]$
@@
Equation (Topological advantage for adsorbate chemisorption on conjugated chains) corresponds to the Head–Gordon–Tully electronic-friction tensor, i.e., the equilibrium, Markov isolated system limit for nuclear motion coupled to a dense manifold of electronic excitations tully1995molecular. We note that more general nonequilibrium open-system treatments can yield non-Markovian friction and additional force contributions beyond a simple Markovian friction tensor chen2018current. Nonetheless, in the assumed equilibrium, isolated system limit with fast decoherence (induced by other molecular or substrate degrees of freedom), the HGT expression has been shown to be consistent with more general open-system formulations of electronic friction dou2017universality, chen2018current.
In our simulations the electronic environment is represented by a finite SSH chain, so the electronic spectrum is discrete at finite $N$. To obtain a continuous spectrum, and hence a well-defined Markovian friction, we broaden each discrete level by a Gaussian of width $\sigma$ following standard practice tully1995molecular, hellsing1984electronic, replacing the Dirac delta functions by normalized Gaussians,
@@
thereby yielding a smooth LDOS $P(\epsilon,R)$ and numerically stable friction coefficients askerka2016role.
In this work, we use an $N$-independent broadening parameter chosen on the order of the thermal smearing scale, $\sigma\sim k_{B}T$ aarons2016perspective. Importantly, our conclusions do not depend on this choice: the friction coefficients are robust with respect to moderate changes in $\sigma$ and with respect to increasing the chain size $N$ (see Supporting Information )
Bulk and boundary chemisorption. In what follows, we examine the chemisorption of a single adsorbate on the SSH chain under various conditions.
To place our parameter choices in the context of realistic 1D conjugated systems, we note that tight-binding descriptions of polyacetylene and polyacetylene-like chains mizes1991tight, meider1997density, wang2019surface, paloheimo1992density are broadly consistent with nearest-neighbor $\pi$-electron hopping scales of a few eV and a moderate bond-alternation (dimerization) contrast. Representative values such as $v=2.5$ eV and $w=3.5$ eV mizes1991tight imply a hopping ratio $w/v$ of order unity with appreciable bond alternation, consistent with the scale of parameters used in the present model.
Because our central conclusions are governed by dimensionless ratios such as the dimerization ratio
@@
the dimensionless LUMO energy $\varepsilon_{0}/v$, and hopping $T/v$, the predicted midgap-state-enabled enhancement of hybridization and electronic friction is robust across the range of $r$ explored here. We therefore expect the reported trends to be transferable.
Our analysis covers the trivial ($r<1$), metallic ($r\simeq 1$), and topological ($r>1$) regimes. We emphasize our primary objective is not to model a specific material parameter set, but to map how the topological phase transition and the appearance of boundary-localized midgap states modulate charge hybridization and electronic friction for an adsorbate across a broad, physically motivated range of $r$.
Throughout this subsection, the molecular geometry is fixed ($R=0$). The adsorbate is therefore characterized by its low-lying level energy $\varepsilon_{0}$ and placement $x_{M}$. We also note that all equilibrium observables reported below are converged with respect to the SSH chain length: increasing the number of unit cells $N$ leaves the electron occupancy essentially unchanged so that the finite chain adopted here represents the system in thermodynamic limit (see Supporting Information for verification).
Figures 2(a,b) report the LUMO occupation $n_{0}(R,x_{M})$ at 0 K and half filling ($\mu=0^{-}$) for adsorption near the edge ($x_{M}=x_{E}$) or at the bulk center ($x_{M}=x_{B}$). Figure 2(a) shows that the trivial insulator exhibits negligible electron sharing at the edge, whereas the topological phase displays strong hybridization with a sharp rise of $n_{0}(R,x_{E})$ upon entering $r>1$. The enhancement is most pronounced when $\varepsilon_{0}$ is nearly resonant with the boundary mode. Conversely, Fig. 2(b) shows that for adsorption in the metallic bulk, significant $n_{0}(R,x_{B})$ occurs only in a narrow neighborhood of the topological transition critical point ($r\approx 1$), where the density of occupied states near $\varepsilon_{0}$ is largest. While the metallic phase maximizes bulk adsorption, boundary chemisorption in the topological phase remains the optimal overall scenario for hybridization under the considered conditions as $\max n_{0}(R,x_{B})<\max n_{0}(R,x_{E})=0.5$. Notably, appreciable adsorbate-edge hybridization persists even when $\varepsilon_{0}\gg 0$.
@@
The robustness of the topological enhancement is highlighted in Fig. 3, which directly compares edge adsorption in the topological phase to bulk adsorption in the metallic phase over varying $\varepsilon_{0}$ and couplings $T_{m,A}$. Across all cases examined, hybridization with metallic bulk states yields a lower $n_{0}$ than coupling to the topological edge mode, despite the larger metallic DOS near $\varepsilon_{0}$. Notably, edge $n_{0}(R,x_{E})$ remains substantial even when $T_{1,A}\ll v,w$.
@@
Discussion. In the topological phase and near resonance ($\varepsilon_{0}\approx 0$), the adsorbate couples predominantly to the nearest edge mode. For $T_{1,A},T_{1,B}\ll v,w$ and $w/v-1\gg 0$, the hybridized edge–LUMO doublet is well isolated from bulk continua, yielding the symmetric occupancy $n_{0}(R,x_{E})\approx 0.5$. In the metallic regime ($r\simeq 1$), $\varepsilon_{0}\approx 0$ is near-degenerate with many occupied/unoccupied extended states and the resulting Fano line shape fano1961effects in the local molecular DOS produces a lower maximal $n_{0}\approx 0.4$. Off resonance, bulk adsorption admits the standard second-order estimate
$n_{0}(R,x_{B})\;\approx\;\sum_{E_{\kappa}^{(0)}<\mu}\frac{\bigl|\boldsymbol{T}_{c}\!\cdot\!\boldsymbol{u}_{\kappa,c}\bigr|^{2}}{\bigl(\varepsilon_{0}-E_{\kappa}^{(0)}\bigr)^{2}},$ (15)
where $E_{\kappa}^{(0)}$ denote bare SSH eigenenergies, $c=N/2$ and $\boldsymbol{T}_{c}$ and $\boldsymbol{u}_{\kappa,c}$ are
$\boldsymbol{T}_{c}\equiv\begin{pmatrix}T_{c,A}\\[2.0pt] T_{c,B}\\[2.0pt] T_{c-1,B}\end{pmatrix},\qquad\boldsymbol{u}_{\kappa,c}\equiv\begin{pmatrix}a_{c,\kappa}\\[2.0pt] b_{c,\kappa}\\[2.0pt] b_{c-1,\kappa}\end{pmatrix}$,
and $a_{j,\kappa},b_{j,\kappa}$ are the amplitudes of eigenstate $\kappa$ on sublattices $A/B$ at cell $j$. The sum runs over occupied states at T=0T\!=\!0 K. Away from criticality the occupied bulk states with non-negligible weight at the molecule are delocalized and detuned from $\varepsilon_{0}$, suppressing $n_{0}(R,x_{B})$. Conversely, when $r\approx 1$, the increased density of states and small denominators in (15) enhance $n_{0}$.